OpenStudy (anonymous):

den shaped like rectangle. perimeter is 432 ft. lenght is 8 tmes as long as width. find length and width.

OpenStudy (anonymous):

let x= width, y= length 2x+2y=432 since length is 8 times longer than width y=8x now substitute 2x+2(8x)=423 2x+16x=432 18x=432 x=24 now y=8x so y=192 therefore width is 24 ft length=192 ft

OpenStudy (anonymous):

Permeter=p=432=2w+2L, but L=8w therefore 432=2w+2(8w) 432= 2w+16w=18w 432/18=24ft=w ........ans L=8w=8(24)=192 ft .....ans